# Bias–variance Tradeoff in Tensor Estimation

### v1

@@ 2 Bias–variance tradeoff in tensor estimation
3 Bias–variance tradeoff in matrix estimation
We consider the model
$Y=X^{*}+Z\in\mathbb{R}^{m\times n},$ (4)
where $X^{*}$ is an unknown matrix with arbitrary rank, and $Z$ is a perturbation (noise) matrix. Even when $\mathrm{rank}(X^{*})$ is not small, it is common to approximate $X^{*}$ by a rank-$r$ estimator, because (i) low-rank decomposition can dramatically increase the computational and memory efficiency, see Remark 3 for more details; and (ii) the approximation bias can be small, so balancing the bias and variance may lead to smaller estimation error. Let
By the Eckart–Young–Mirsky theorem (Eckart and Young,, 1936), $X_{(r)}^{*}$ is the best rank-$r$ approximation to $X^{*}$ in Frobenius norm in the sense that
Suppose we observe $Y$ instead of $X^{*}$, and write the SVD of $Y$ as
Theorem 2.
Under (4), let $Y_{(r)}$ be defined as in (5). Then
@@ 3 Bias–variance tradeoff in matrix estimation
Remark 3.
$t=V_{(r)}^{\top}v\in\mathbb{R}^{r}\ (\text{cost }\approx 2nr),\quad s=\Sigma_{(r)}t\in\mathbb{R}^{r}\ (\text{cost }\approx r),\quad w=U_{(r)}s\in\mathbb{R}^{m}\quad(\text{cost }\approx 2mr).$
Beyond computational savings, the SVD factorization also reduces storage costs. Storing the full matrix $Y$ requires $O(mn)$ memory, whereas storing the factors $U_{(r)}\in\mathbb{R}^{m\times r}$, $\Sigma_{(r)}\in\mathbb{R}^{r\times r}$, and $V_{(r)}\in\mathbb{R}^{n\times r}$ requires only $O(mr+nr)$ scalars.
In the following, we illustrate the application of Theorem 2 in three matrix estimation settings.
Corollary 3.
Suppose (4) holds and that the entries of $Z$ are independent sub-Gaussian random variables with
$\|Z_{ij}\|_{\psi_{2}}\leq\kappa.$
$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\big(\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\kappa\sqrt{r(m+n)}\big).$
Proof of Corollary 3.
From Vershynin, (2018) section 4.4.2, it follows that $\|Z\|\leq C_{1}\kappa\sqrt{m+n}$ with probability at least $1-\exp(-(m+n))$ . The desired result follows immediately from Theorem 2. ∎
Corollary 4.
$\|v^{\top}Zw\|_{\psi_{2}}\leq\kappa.$
$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\big(\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\kappa\sqrt{r(m+n)}\big).$
@@
Corollary 5.
$Y=\frac{1}{N}\sum_{k=1}^{N}Z_{k}Z_{k}^{\top},$
$\|Y_{(r)}-X^{*}\|_{\mathrm{F}}\leq C\Bigl(\|X_{(r)}^{*}-X^{*}\|_{\mathrm{F}}+\kappa^{2}\sqrt{r}\Bigl[\sqrt{\frac{n}{N}}+\frac{n}{N}\Bigr]\Bigr).$
@@
Let $\widetilde{Z}:=Y-X^{*}.$ By Remark 4.7.3 of Vershynin, (2018), with probability at least $1-\exp(-cn)$ one has $\|\widetilde{Z}\|\;\leq\;C_{1}\kappa^{2}\Bigl[\sqrt{\frac{n}{N}}+\frac{n}{N}\Bigr].$ Applying Theorem 2 with $Z$ replaced by $\widetilde{Z}$ yields the desired bound. ∎
4 Numerical experiments
In many applied settings, selecting the single best rank $r$ is inherently difficult in finite samples with unknown latent structure. Our analysis instead provides guarantees for singular-value decompositions in both the matrix (SVD) and tensor (HOSVD) regimes across a broad range of $r$. Consequently, rather than hinging on a brittle rank-selection heuristic, we empirically demonstrate that both truncated SVD and one-step HOSVD (Algorithm 1) are robust whenever ranks are chosen neither too small (underfitting/bias) nor too large (overfitting/variance). This is consistent with the bias–variance tradeoff formalized in Theorems 2 and 1. Moreover, in practice the choice of $r$, is frequently driven by diverse aims other than merely minimizing mean-squared error, including visualization, interpretability, or downstream decision tasks. This is reflected in widespread, long-standing use across biology/genomics (Price et al.,, 2006), economics (Filmer and Pritchett,, 2001), and computer vision (Turk and Pentland,, 1991).
We evaluate these predictions on (i) a real 3D brain-MRI dataset with controlled additive noise (IXI; T1-weighted volumes (IXI,, 2002)), and (ii) simulated data for both matrices and third–order tensors.
4.1 Real data: 3D brain MRI with controlled noise
We evaluate on a T1-weighted brain MRI volume drawn from the IXI dataset (IXI,, 2002). The selected volume is a 3D tensor of shape $256\times 256\times 150$. To probe low-rank behavior in a realistic preprocessing pipeline, the volume is prepared in two wavelet-smoothed variants using standard orthonormal families—(a) Daubechies-6 (db6) and (b) Symlet-8 (sym8) (Daubechies,, 1988, 1992).
We implement separable 3D multilevel decompositions (wavedecn) with periodic boundary handling, and apply subband-wise soft-thresholding via BayesShrink (Chang et al.,, 2000; Donoho and Johnstone,, 1994). The literature supports wavelet shrinkage as an effective denoising/smoothing step for MR magnitude images (Nowak,, 1999; Wood and Johnson,, 1999).
Noise model and ranks. For our prepared volume $X$ and each noise level $\lambda\in\{0.01,0.05,0.1,0.2\}$ we generate
which keeps the per-entry SNR comparable across images and variants. We compute one-step HOSVD reconstructions (Algorithm 1) $\widetilde{X}$ at Tucker rank $(r,r,r)$ for $r\in\{50,65,80\}.$
Findings. Figure 1 shows representative mid-sagittal slices from the reconstructions of the wavelet-smoothed variants of single volume. Qualitatively, the slices exhibit the expected progression: very small ranks $r$ lead to underfitting with over-smoothed, low-rank artifacts, while an intermediate range of $r$ preserves cortical detail and simultaneously suppresses background fluctuations.
Figure 1: Mid-sagittal MRI slice of the reconstructed volume at selected ranks and noise levels $\lambda$ for the db6 variant (left), and for the sym8 variant (right).
4.2 Synthetic data: Tensors
We generate third-order tensors $X^{*}\in\mathbb{R}^{p\times p\times p}$ in the Tucker form
$X^{*}\;=\;\mathcal{G}\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3},\qquad U_{1},U_{2},U_{3}\in\mathbb{R}^{p\times s},\ U_{1}^{\top}U_{1}=U_{2}^{\top}U_{2}=U_{3}^{\top}U_{3}=I_{s},$
where $U_{1},U_{2},U_{3}$ are drawn independently with orthonormal columns. The core $\mathcal{G}\in\mathbb{R}^{s\times s\times s}$ is diagonal along its main fiber with entries $\{\gamma_{i}\}_{i=1}^{s}$ decaying exponentially, i.e. $\gamma_{i}=\beta^{i},\,\beta=0.8$. We observe $Y=X^{*}+Z$ with i.i.d. Gaussian noise $Z_{ijk}\sim\mathcal{N}(0,1)$ entry-wise. The SNR is controlled by scaling the signal such that
$\|X^{*}\|_{\mathrm{F}}\;=\;\lambda\,\sqrt{p^{3}},\qquad\lambda\in\{10,50\}.$
Dimensions and evaluated ranks. We vary $p\in\{20,50,75,100\}$ and for each $(p,s)$, we evaluate two ranks $r$ as listed below:
$p$ $s$ $r$ (target)
20 15 10, 12
50 25 10, 15
75 20 10, 15
100 80 30, 40
Table 1: Synthetic tensor grid: dimensions and evaluated Tucker ranks.
Each configuration $(p,s,\lambda,r)$ is repeated $R=50$ times with freshly drawn $(U_{1},U_{2},U_{3})$ and noise. Implementations use exact batched SVD on GPU via TensorLy (PyTorch backend) (Kossaifi et al.,, 2019). We report the sample mean and standard error of the relative Frobenius error
$\mathrm{RelErr}(\widetilde{X};X^{*})\;=\;\frac{\|\widetilde{X}-X^{*}\|_{\mathrm{F}}}{\|X^{*}\|_{\mathrm{F}}}\,$,
4.3 Synthetic data: Matrices
$X^{*}\;=\;U\,\Sigma\,V^{\top},\qquad U\in\mathbb{R}^{m\times n},\ \ V\in\mathbb{R}^{n\times n},\ \ U^{\top}U=V^{\top}V=I_{n},$
where $\Sigma=\mathrm{diag}(\sigma_{1},\ldots,\sigma_{n})$ has exponentially decaying singular values, i.e. $\sigma_{i}=\beta^{\,i}$ with $\beta=0.8$. To mirror the tensor ambient sizes while probing rectangularity, we set $m=5p$ and $n=s$ according to the tensor grid in Table 1. We observe $Y=X^{\star}+Z$, where $Z_{ij}\sim\mathcal{N}(0,1)$ is i.i.d. Gaussian noise entry-wise. The SNR is controlled by scaling the signal such that
$\|X^{*}\|_{\mathrm{F}}\;=\;\lambda\,\sqrt{mn},\qquad\lambda\in\{10,50\}.$
Table 2: Results for matrix (left) and tensor (right).
Matrix
$\lambda$ $m$ $n$ $r$ Mean (SE)
10 100 15 10 0.1178 (0.00125)
12 0.0906 (0.00164)
250 25 10 0.1254 (0.00044)
15 0.0868 (0.00077)
375 20 10 0.1279 (0.00045)
15 0.0927 (0.00076)
500 80 30 0.0698 (0.00032)
40 0.0795 (0.00035)
50 100 15 10 0.0844 (0.00007)
12 0.0181 (0.00037)
250 25 10 0.1079 (0.00002)
15 0.0378 (0.00009)
375 20 10 0.1068 (0.00002)
15 0.0349 (0.00008)
500 80 30 0.0134 (0.00008)
40 0.0156 (0.00006)
Tensor
$\lambda$ $p$ $s$ $r$ Mean (SE)
10 20 15 10 0.1092 (0.00031)
12 0.0776 (0.00058)
50 25 10 0.1081 (0.00002)
15 0.0402 (0.00012)
75 20 10 0.1081 (0.00002)
15 0.0353 (0.00004)
100 80 30 0.0204 (0.00007)
40 0.0294 (0.00007)
50 20 15 10 0.1018 (0.00001)
12 0.0599 (0.00002)
50 25 10 0.1073 (0.00000)
15 0.0352 (0.00000)
75 20 10 0.1067 (0.00000)
15 0.0333 (0.00000)
100 80 30 0.0039 (0.00001)
40 0.0058 (0.00001)
5 Discussion
Future work.
(i) Beyond i.i.d. noise and full observations. Extending the theory to heteroskedastic or correlated perturbations (e.g. spatially correlated fields, Rician-like MRI noise) and to incomplete observations (tensor completion, masked entries) is natural. We expect the variance term to inherit problem-dependent effective dimensions (e.g. leverage scores or sampling densities), while the bias term remains $\xi_{(r_{1},r_{2},r_{3})}$. A key technical step is replacing isotropic concentration with mode-wise covariance-aware bounds for $\mathcal{M}_{k}(Z)$ and their projections. (ii) Higher orders and higher ranks. The proof strategy appears to scale to $d$th-order tensors with the clean generalization
Without loss of generality, if $W\sim\mathrm{subGaussian}(0,\kappa^{2})$, we assume throughout, unless otherwise specified in the appendix, that $\mathbb{E}[W^{2}]=\kappa^{2}$.
Appendix A Proof of Theorem 1

### v3

@@ 3 Optimal bias–variance tradeoff in tensor estimation
4 Numerical experiments
We numerically demonstrate the bias-variance tradeoff developed in our theory. We begin by describing a practical approach of selecting the rank parameters used in both SVD and one-step HOSVD (Algorithm 1).
Remark 4 (Practical choice of target ranks).
There are several well-known approaches for selecting the target Tucker ranks. To select the rank parameters in a data-driven manner, we recommend following the discussion in Ozdemir et al. (2017) and adopting an adaptive thresholding scheme. More precisely, one selects $r_{k}$ such that $\sigma_{r_{k}}^{2}(\mathcal{M}_{k}(Y))\geq\tau$, where $\tau>0$ is prescribed threshold. A simple and effective choice is $\tau=\alpha\|Y\|_{\mathrm{F}}^{2}$ with $\alpha=0.01$. Under this rule, all singular values of $\mathcal{M}_{k}(Y)$ up to the largest index $j$ satisfying $\sigma_{j}^{2}(\mathcal{M}_{k}(Y))\geq\tau$ are retained, and the remainder are discarded.
In many applied settings, however, selecting the single best rank $r$ is inherently difficult in finite samples, particularly with unknown latent structure. Our analysis instead provides guarantees for singular-value decompositions in both the matrix and tensor regimes across a broad range of $r$. Consequently, rather than relying on a specific rank-selection heuristic, we empirically demonstrate that both truncated SVD and one-step HOSVD are robust provided that the chosen ranks are neither too small (underfitting) nor too large (overfitting). This behavior is consistent with the bias–variance tradeoff formalized in Theorems 1 and 3. Moreover, in practice the choice of $r$ is often guided by considerations beyond minimizing mean-squared error, such as visualization, interpretability, or performance on downstream tasks. This perspective aligns with the widespread and long standing use of low-rank methods across diverse application domains, biology and genomics (Price et al., 2006), economics (Filmer and Pritchett, 2001), and computer vision (Turk and Pentland, 1991).
We evaluate the bias-variance tradeoff on (i) a real 3D brain-MRI dataset with controlled additive noise (Group, 2002, T1-weighted volumes), and (ii) simulated data in both matrix and third–order tensor settings.
4.1 Real data: 3D brain MRI with controlled noise
We evaluate on five T1-weighted brain MRI volumes drawn from the IXI dataset (Group, 2002). The selected volumes are third-order tensors of shape $256\times 256\times 150$.
For each volume, we consider the original tensor and two wavelet-smoothed tensors obtained using orthonormal wavelet families Daubechies-6 (db6) and Symlet-8 (sym8). Wavelet smoothing is performed using separable 3D multilevel decompositions (wavedecn) followed by subband-wise soft-thresholding with BayesShrink (Donoho and Johnstone, 1994; Chang et al., 2000). This approach is well supported in the MRI literature as an effective smoothing strategy for MRI volumes (Nowak, 1999; Wood and Johnson, 1999).
Noise model and ranks. For each noise level $\lambda\in\{0.2,0.3,0.4,0.5,0.6,0.7\}$ and each volume $X^{\ast}$, we generate
which keeps the per-entry SNR comparable across images and variants. We compute one-step HOSVD reconstructions (Algorithm 1) $\widetilde{X}$ at Tucker rank $(r,r,r)$ for $r\in\{35,50,60,65,70,80,90,95\}.$
Findings. Figure 1 displays mid-sagittal slices from reconstructions of a selected volume, comparing results obtained from inputs smoothed with db6 and sym8 wavelets. At intermediate ranks, the images sharpen and cortical detail appears while background fluctuations remain suppressed. We provide selected quantitative results in Table 1, which show this expected bias-variance tradeoff progression.
Figure 1: Mid-sagittal MRI slices reconstructed with Algorithm 1 at Tucker ranks $(r,r,r)$ across noise levels $\lambda$ for a selected volume. Inputs smoothed with db6 (left) and sym8 (right) wavelets.
Table 1: Mean and standard deviation (SD) of the relative reconstruction error with 55 MRI volumnes. Results are shown across varying scaling parameters $\lambda$ and ranks $r$.
$\lambda$ $r$ original db6 sym8
0.4 50 0.1753 (0.0330) 0.1747 (0.0537) 0.1681 (0.0466)
65 0.1553 (0.0268) 0.1561 (0.0435) 0.1501 (0.0373)
80 0.1490 (0.0183) 0.1499 (0.0291) 0.1451 (0.0245)
95 0.1550 (0.0102) 0.1558 (0.0159) 0.1525 (0.0128)
0.5 50 0.1797 (0.0322) 0.1793 (0.0521) 0.1728 (0.0451)
65 0.1655 (0.0248) 0.1668 (0.0400) 0.1610 (0.0341)
80 0.1677 (0.0154) 0.1691 (0.0244) 0.1647 (0.0201)
95 0.1840 (0.0073) 0.1847 (0.0116) 0.1819 (0.0089)
0.6 50 0.1851 (0.0311) 0.1850 (0.0503) 0.1787 (0.0434)
65 0.1776 (0.0224) 0.1793 (0.0362) 0.1738 (0.0307)
80 0.1890 (0.0124) 0.1904 (0.0200) 0.1866 (0.0162)
95 0.2150 (0.0049) 0.2154 (0.0085) 0.2132 (0.0062)
4.2 Synthetic data: Tensors
We generate third-order tensors $X^{\ast}\in\mathbb{R}^{p\times p\times p}$ in the Tucker form
$X^{\ast}=\mathcal{G}\times_{1}U_{1}\times_{2}U_{2}\times_{3}U_{3},\qquad U_{1},U_{2},U_{3}\in\mathbb{R}^{p\times p},\ U_{1}^{\top}U_{1}=U_{2}^{\top}U_{2}=U_{3}^{\top}U_{3}=I_{p},$
where $U_{1},U_{2},U_{3}$ are drawn independently with orthonormal columns. The core tensor $\mathcal{G}\in\mathbb{R}^{p\times p\times p}$ is defined entrywise by
$\mathcal{G}_{ijk}=\gamma_{ijk}\beta^{i+j+k},\qquad\beta=0.8,$
where the signs $\gamma_{ijk}$ are independent with $\mathbb{P}(\gamma_{ijk}=-1)=1/3,\mathbb{P}(\gamma_{ijk}=1)=2/3$. We observe $Y=X^{\ast}+Z$ with i.i.d. Gaussian noise $Z_{ijk}\sim\mathcal{N}(0,1)$ entry-wise. The SNR is controlled by scaling the signal such that
$\|X^{\ast}\|_{\mathrm{F}}=\lambda\sqrt{p^{3}},\qquad\lambda\in\{10,50\}.$
Dimensions and evaluated ranks. We vary $p\in\{20,50,75,100\}$ and for each $p$, we evaluate two ranks $r$ as outlined in Table 2.
Each configuration $(p,\lambda,r)$ is repeated $R=50$ times with freshly drawn $(U_{1},U_{2},U_{3})$ and noise. Implementations use exact batched SVD on GPU via TensorLy (PyTorch backend) (Kossaifi et al., 2019). We report the sample mean and standard deviation of the relative Frobenius error
$\mathrm{RelErr}(\widetilde{X};X^{\ast})=\frac{\|\widetilde{X}-X^{\ast}\|_{\mathrm{F}}}{\|X^{\ast}\|_{\mathrm{F}}}\,$,
Bias–variance curve. To empirically quantify the contributions of bias and variance to the error $\|\widetilde{X}-X^{\ast}\|_{\mathrm{F}}$, we consider out synthetic tensor model with $p=100$. For each candidate Tucker rank $r\in\{12,15,18,\ldots,84,87\}$, we run $50$ independent repetitions. The bias component is obtained as the error $\|\widetilde{X}-X^{\ast}\|_{\mathrm{F}}$, where $\widetilde{X}$ denotes the output of Algorithm 1 when applied directly to the noiseless tensor $X^{\ast}$. The variance component is obtained as
$\max\left\{\sqrt{r}\left\|\mathcal{M}_{1}(X^{\ast})\cdot\left\{U_{2}^{(1)}\otimes U_{3}^{(1)}\right\}\right\|,\,\sqrt{r}\left\|\mathcal{M}_{2}(X^{\ast})\cdot\left\{U_{1}^{(1)}\otimes U_{3}^{(1)}\right\}\right\|,\,\sqrt{r}\left\|\mathcal{M}_{3}(X^{\ast})\cdot\left\{U_{1}^{(1)}\otimes U_{2}^{(1)}\right\}\right\|,\,\left\|X^{\ast}\times_{1}U_{1}^{(1)}\times_{2}U_{2}^{(1)}\times_{3}U_{3}^{(1)}\right\|_{\mathrm{F}}\right\}.$
Figure 2 reports the resulting mean bias and variance curves as functions of the target rank: as the target rank increases, the bias term decreases while the variance term grows, leading to an intermediate rank that minimizes the overall error.
Figure 2: Mean bias and variance components of the error as functions of the target rank in the synthetic experiment with $p=100$. The left panel represent the curve with scaling parameter $\lambda=10$ and the right panel $\lambda=50$.
4.3 Synthetic data: Matrices
$X^{\ast}=U\,\Sigma\,V^{\top},\qquad U\in\mathbb{R}^{m\times n},\ \ V\in\mathbb{R}^{n\times n},\ \ U^{\top}U=V^{\top}V=I_{n},$
where $\Sigma=\mathrm{diag}(\sigma_{1},\ldots,\sigma_{n})$ has exponentially decaying singular values, i.e. $\sigma_{i}=\beta^{\,i}$ with $\beta=0.8$. We choose $m$ and $n$ as outlined in Table 2. We observe $Y=X^{\ast}+Z$, where $Z_{ij}\sim\mathcal{N}(0,1)$ is i.i.d. Gaussian noise entry-wise. The SNR is controlled by scaling the signal such that
$\|X^{\ast}\|_{\mathrm{F}}=\lambda\sqrt{mn},\qquad\lambda\in\{10,50\}.$
Table 2: Mean and standard deviation (SD) of the relative error (RelErr) for matrices (left) and tensors (right). Results are shown across varying scaling parameters $\lambda$, dimensions ($m\times n$ for matrices; $p\times p\times p$ for tensors), and ranks $r$. For each configuration, results are averaged over $R=50$ Monte Carlo simulations.
Matrix
$\lambda$ $m$ $n$ $r$ Mean (SD)
10 100 15 10 0.1178 (0.00125)
12 0.0906 (0.00164)
250 25 10 0.1254 (0.00044)
15 0.0868 (0.00077)
375 20 10 0.1279 (0.00045)
15 0.0927 (0.00076)
500 80 30 0.0698 (0.00032)
40 0.0795 (0.00035)
50 100 15 10 0.0844 (0.00007)
12 0.0181 (0.00037)
250 25 10 0.1079 (0.00002)
15 0.0378 (0.00009)
375 20 10 0.1068 (0.00002)
15 0.0349 (0.00008)
500 80 30 0.0134 (0.00008)
40 0.0156 (0.00006)
Tensor
$\lambda$ $p$ $r$ Mean (SD)
10 20 10 0.1420 (0.00350)
12 0.0977 (0.00208)
50 10 0.1370 (0.00347)
15 0.0459 (0.00102)
75 10 0.1367 (0.00348)
15 0.0427 (0.00113)
100 30 0.0203 (0.00008)
40 0.0294 (0.00009)
50 20 10 0.1361 (0.00381)
12 0.0840 (0.00207)
50 10 0.1362 (0.00359)
15 0.0410 (0.00109)
75 10 0.1368 (0.00374)
15 0.0412 (0.00098)
100 30 0.0039 (0.00002)
40 0.0058 (0.00002)
5 Discussion
Future work. (i) Beyond i.i.d. noise and full observations. Extending the theory to heteroskedastic or correlated perturbations (e.g. spatially correlated fields, Rician-like MRI noise) and to incomplete observations (tensor completion, masked entries) is natural. We expect the variance term to inherit problem-dependent effective dimensions (e.g. leverage scores or sampling densities), while the bias term remains $\xi_{(r_{1},r_{2},r_{3})}$. A key technical step is replacing isotropic concentration with mode-wise covariance-aware bounds for $\mathcal{M}_{k}(Z)$ and their projections. (ii) Higher orders and higher ranks. The proof strategy appears to scale to $d$th-order tensors with the clean generalization
Throughout this appendix we work with centered sub-Gaussian random variables with parameter $\kappa^{2}$. we assume, unless otherwise specified, that each such variable $X$ satisfies $\mathbb{E}X^{2}=\kappa^{2}$.
Appendix A Proofs in Section 2
@@
Proof of Theorem 2.
Without loss of generality, suppose that $m\leq n$ and $r\leq m/2$. Therefore it suffices to show that
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{M}_{r}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c_{1}\Bigl\{\kappa^{2}nr+\xi^{2}\Bigr\}.$ (13)
Let $r^{\prime}=\lfloor\xi^{2}/(n\kappa^{2})\rfloor$. Note that since $\xi^{2}\leq(m-r)n\kappa^{2}$, it follows that
$r^{\prime}+r\leq m.$
Step 1. Consider the set of matrices
$\mathcal{B}=\{B\in\mathbb{R}^{m\times n}:\|B\|_{\infty}\leq\kappa^{2},B_{i,j}=0\text{ if }i>r+r^{\prime}\}.$
In this step, we show that
$\mathcal{B}\subset\mathcal{M}^{\xi}_{r}.$ (14)
To see this, let $B\in\mathcal{B}$, and denote $B^{\prime}$ to be the matrix such that the first $r$ rows of $B^{\prime}$ equals the first $r$ rows of $B$, and the rest of (mr)(m-r) rows of $B^{\prime}$ are all 0. Then since $B^{\prime}$ has at most $r$ rows that are nonzero, $\operatorname*{rank}(B^{\prime})\leq r$ and
$\|B-B^{\prime}\|_{\mathrm{F}}^{2}\leq r^{\prime}n\|B\|_{\infty}=r^{\prime}n\kappa^{2}\leq\xi^{2}.$
Therefore $\mathcal{B}\subset\mathcal{M}^{\xi}_{r}$. So
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{M}_{r}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\ \geq\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}.$ (15)
Step 2. Note that matrices in $\mathcal{B}$ are having at most $(r+r^{\prime})n$ nonzero entries, the set of matrices $\mathcal{B}$ can be viewed as vectors in $\mathbb{R}^{d}$, with $d=(r+r^{\prime})n$. Denote
$\mathcal{V}=\{V\in\mathbb{R}^{d},\|V\|_{\infty}\leq\kappa\}.$
Then
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}=\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}.$
Since by Lemma 19, $\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}\geq c_{2}\kappa^{2}d,$ it follows that
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{M}_{r}^{\xi}}\mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\geq\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{B}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}=\inf_{\widehat{V}}\sup_{V\in\mathcal{V}}\mathbb{E}\|\widehat{V}-V\|^{2}\geq c_{2}\kappa^{2}d=c_{2}\kappa^{2}(r+r^{\prime})n.$
Step 3. There are two cases. (a) Suppose $\xi^{2}\geq n\kappa^{2}$, then $r^{\prime}=\lfloor\xi^{2}/(n\kappa^{2})\rfloor\geq 1$ and therefore
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{M}_{r}^{\xi}}\mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\geq c_{2}\kappa^{2}(r+r^{\prime})n\geq c_{2}rn+c_{3}\xi^{2}.$
(b) Suppose $\xi^{2}<n\kappa^{2}$. Then $r^{\prime}=0$
$\inf_{\widehat{X}}\sup_{X^{*}\in\mathcal{M}_{r}^{\xi}}\ \mathbb{E}\bigl\|\widehat{X}-X^{*}\bigr\|_{\mathrm{F}}^{2}\$$\geq\ c_{2}\kappa^{2}rn\geq\frac{c_{2}}{2}(\kappa^{2}rn+\xi^{2}),$ (16)
where the first inequality follows from Step 1 and Step 2, the second inequality follows from $\xi^{2}<n\kappa^{2}$. ∎
@@
Let $\widetilde{Z}=Y-X^{*}.$ By Remark 4.7.3 of Vershynin (2018), with probability at least $1-\exp(-cn)$ one has $\|\widetilde{Z}\|\;\leq\;C_{1}\kappa^{2}\Bigl[\sqrt{\frac{n}{N}}+\frac{n}{N}\Bigr].$ Applying Theorem 1 with $Z$ replaced by $\widetilde{Z}$ yields the desired bound. ∎
Appendix B Proofs in Section 3